40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 16:00:01 2019, gautam, Update, CDS, c1iscaux Supermicro setup IMG_7769.JPG
    Reply  Wed Jul 17 21:22:41 2019, gautam, Update, CDS, CM board Latch Enable subtlety 
       Reply  Sun Jul 21 12:55:38 2019, gautam, Update, CDS, CM board Latch Enable test script LatchLogic.pdfLatchLogicTest.png
          Reply  Fri Aug 9 08:59:04 2019, gautam, Update, CDS, Prep for install of c1iscaux 
             Reply  Sun Aug 11 11:47:42 2019, gautam, Update, CDS, Bench test of c1iscaux 
                Reply  Mon Aug 12 17:36:04 2019, gautam, Update, CDS, More bench test of c1iscaux iscauxCheclist.pdf
                   Reply  Mon Aug 12 21:25:19 2019, Koji, Update, CDS, More bench test of c1iscaux 
                   Reply  Tue Aug 13 08:07:09 2019, gautam, Update, CDS, P1--->P2 
                      Reply  Tue Aug 13 14:36:17 2019, gautam, Update, CDS, P1--->P2 
                   Reply  Fri Aug 16 16:40:04 2019, gautam, Update, CDS, 1Y3 work newLook1Y3.JPGIMG_7803.JPGc1lsc_crashed.png
                      Reply  Sat Aug 17 16:49:23 2019, gautam, Update, CDS, More 1Y3 work Screen_Shot_2019-08-17_at_3.00.57_PM.pngScreen_Shot_2019-08-17_at_3.12.23_PM.pngIMG_7804.JPGScreenshot_from_2019-08-17_17-04-47.png
                         Reply  Mon Aug 19 14:36:21 2019, gautam, Update, CDS, c1iscaux remaining work caseForSmallerFootprint.pdf
                            Reply  Tue Aug 20 19:05:24 2019, Koji, Update, CDS, MC1 (and MC3) troubleshoot Screenshot_from_2019-08-20_17-26-01.pngScreenshot_from_2019-08-20_17-43-03.png
                               Reply  Thu Aug 22 12:54:06 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 7x
                                  Reply  Thu Aug 22 20:56:51 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 
                            Reply  Fri Aug 23 18:46:17 2019, Jon, Update, CDS, c1iscaux remaining work 
                               Reply  Sun Aug 25 14:18:08 2019, gautam, Update, CDS, c1iscaux remaining work Screen_Shot_2019-08-25_at_10.38.37_PM.png
    Reply  Thu Jul 18 10:46:04 2019, gautam, Update, CDS, Database files made 
       Reply  Thu Jul 18 19:58:56 2019, gautam, Update, CDS, Work on Acromag chassis IMG_7771.JPGIMG_7770.JPGIMG_7772.JPG
          Reply  Fri Jul 19 19:44:03 2019, gautam, Update, CDS, Database file test Whitening.png
Message ID: 14852     Entry time: Thu Aug 22 12:54:06 2019     In reply to: 14851     Reply to this: 14853
Author: Koji 
Type: Update 
Category: CDS 
Subject: MC1 glitch removed (for now) and IMC locking recovered 

I have checked the MC1 satellite box and made a bunch of changes. For now, the glitches coming from the satellite box is gone. I quickly tested the MC1 damping and the IMC locking. The IMC was locked as usual. I still have some cleaning up but will work on them today and tomorrow.


Attachment 1: Result

The noise level of the satellite box was tested with the suspension simulator (i.e., five pair of the LED and PD in a plastic box).

Each plot shows the ASD of the sensor outputs 1) before the modification, 2) after the change, and 3) with the satellite box disconnected (i.e., the noise from the PD whitening filter in the SUS rack).

Before the modification, these five signals showed significant (~0.9) correlation each other, indicating that the noise source is common. After the modification, the spectra are lowered down to the noise level of the whitening filters, and there is no correlation observed anymore. EXCEPT FOR the LR sensor: It seems that the LR has additional noise issue somewhere in the downstream. This is a separate issue.


Attachment 2: Photo of the satellite box before the modification

The thermal environment in the box is terrible. They are too hot to touch. You can see that the flat ribbon cable was burned. The amps, buffers, and regulators generate much heat.

Attachment 3: Where the board was modified

- (upper left corner) Every time I touched C51, the diode output went to zero. So C51 was replaced with WIMA 10uF (50V) cap.

- (lower left area) I found a clear indication of the glitch coming from the PD bias path (U3C). So I first replaced another 10uF (C50) with WIMA 10uF (50V). This did not change the glitch. So I replaced U3 (LT1125). This U3 had unused opamp which had railed to the supply voltage. Pins 14 and 15 of U3 were shorted to ground.

- (lower right corner) Similarly to U3, U6 also had two opamps which are railed due to no termination. U6 was replaced, and Pins 11, 12, 14, and 15 were shorted to ground.

- (middle right) During the course of the search, I suspected that the LR glitch comes from U5. So U5 was replaced to the new chip, but this had no effect.


Attachment 4: Thermal degradation of the internal ribbon cable

Because of the heat, the internal ribbon cable lost the flexibility. The cable is cracked and brittle. It now exposes  some wires. This needs to be replaced. I'll work on this later this week.

Attachment 5: Thermal degradation of the board

Because of the excessive heat for those 20years, the bond between the board and the patten were degraded. In conjunction with extremely thin wire pattern, desoldering of the components (particularly LT1125s) was very difficult. I'd want to throw away this board right now if it were possible...

Attachment 6: Shorting the unused opamps

This shows how the pieces of wires were soldered to ground vias to short the unused opamps.


Attachment 7: Comparison of the noise level with the sus simulator and the actual MC1 motion

After the satellite box fix, the sensor outputs were measured with the suspension connected. This shows that the suspension is moving much more than the noise level around 1Hz. However, at the microseismic frequency there is also most no mergin. Considering the use of the adaptive feedforward, we need to lower the noise of the satellite box as well as the noise of the whitening filters.

=> Use better chips (no LT1125, no current buffers), use low noise resistors, better thermal environment.


Attachment 1: satellite_box.pdf  1.125 MB  Uploaded Thu Aug 22 13:54:20 2019  | Hide | Hide all
satellite_box.pdf
Attachment 2: before.jpg  2.382 MB  Uploaded Thu Aug 22 13:55:44 2019  | Hide | Hide all
before.jpg
Attachment 3: after.jpg  3.272 MB  Uploaded Thu Aug 22 13:56:14 2019  | Hide | Hide all
after.jpg
Attachment 4: P_20190821_194035.jpg  2.486 MB  Uploaded Thu Aug 22 13:56:26 2019  | Hide | Hide all
P_20190821_194035.jpg
Attachment 5: P_20190821_174240.jpg  1.160 MB  Uploaded Thu Aug 22 13:56:36 2019  | Hide | Hide all
P_20190821_174240.jpg
Attachment 6: P_20190821_194013.jpg  2.185 MB  Uploaded Thu Aug 22 13:56:36 2019  | Hide | Hide all
P_20190821_194013.jpg
Attachment 7: comparison_satellite_box.pdf  1.152 MB  Uploaded Thu Aug 22 13:57:32 2019  | Hide | Hide all
comparison_satellite_box.pdf
ELOG V3.1.3-